# Detecting Toxic Flow

### v1

@@ 7.3 The added value of having one model per client
For PULSE, with the exception of Client 1, the performance of one model per client is significantly lower than that of the universal model. Indeed, the extra data is more valuable than a model per client. The universal model employs features that are built using identity of the client, e.g., inventory and cash of client. Thus, it is more advantageous to have one model for all clients and benefit from more data points than one model per client at the expense of having fewer data points to train the individual models. This is the case for PULSE and RF, whereas the results for LogR and MLE are not conclusive. We also find that there is no statistical advantage to consider client-specific features, i.e., cash, inventory, and recent activity of clients. More precisely, accuracies remain the same when we run the models without cash, inventory, and recent activity of clients.
Lastly, we study the added value of employing data from more clients to build the models. Figure 18 shows the daily AUC distribution by toxic horizon when the models are trained with the top one hundred clients by number of trades. We do not consider client features and we use a single global model as opposed to one model per client.
Figure 18: Left panel shows the kernel density estimate plot of daily AUC by toxic horizon. Right panel shows the daily AUC of all models for a toxicity horizon of 30s. The dashed vertical lines correspond to the mean daily value. EUR/USD currency pair over the period 1 August 2022 to 21 October 2022.
The mean daily AUC for all models performs better in Figure 18 (one global model, no features with client information, and top one hundred clients) than that in Figure 8 (one model per client, features with client information, and top six clients). In particular, the outperformance of PULSE (in AUC) ranges from 1% to 9% across toxicity horizons. Thus, in our dataset, having more clients – hence more data – is advantageous. Note that PULSE is the best model.
8 Conclusions
We employed machine learning and statistical methods to detect toxic flow. We also developed a new method, which we call PULSE, for brokers to assess the toxicity of each trade they do with their clients. Out-of-sample AUC of PULSE is high, stable, and outperforms the other methods we considered. We proposed a broker’s strategy that uses these predictions to decide which trades are internalised and which are externalised by the broker. The mean PnL of the internalise-externalise strategy we obtain with PULSE is the highest (when compared with the benchmarks) and it is robust to model parameter choices. Future research will consider a hierarchical version of the problem, where there is a structure for toxicity common to all traders. PULSE can also be used in other areas of financial mathematics, such as in the prediction of fill-rate probabilities, and in multi-armed bandit problems for trading; see, e.g., Arroyo et al., (2023) and Cartea et al., (2023).
Appendix A PULSE derivation and proofs
In this section, we present the derivation and proofs for PULSE. Proposition 1 derives the general fixed point equations. Given that these are expensive to compute, we present additional results to obtain the computationally efficient form of the theorem. We then prove Theorem 2 in A.1.
Proposition 1.
(Modified R-VGA for PULSE) Suppose $\log p(y_{t}\,|\,{\bf z},{\bf w};{\bf x}_{t})$ is differentiable with respect to to $({\bf z},{\bf w})$ and the observations $\{y_{t}\}_{t=1}^{T}$ are conditionally independent over $({\bf z},{\bf w})$. Given Gaussian prior distributions $\phi_{0}$, $\varphi_{0}$ for ${\bf w}$ and ${\bf z}$ respectively, the variational posterior distributions at time $t\in\{1,\ldots,T\}$ that solve (3.6) satisfy the fixed-point equations
${\boldsymbol{\nu}}_{t}$ $={\boldsymbol{\nu}}_{t-1}-{\boldsymbol{\Sigma}}_{t-1}\nabla_{{\boldsymbol{\nu}}}\,{\mathbb{E}}_{{\phi_{t}\left({\bf w}\right)},{\varphi_{t}\left({\bf z}\right)}}[\log p(y_{t}\,|\,{\bf z},{\bf w};{\bf x}_{t})]\,,$ (A.1)
${\boldsymbol{\mu}}_{t}$ $={\boldsymbol{\mu}}_{t-1}-{\boldsymbol{\Gamma}}_{t-1}\nabla_{{\boldsymbol{\mu}}}\,{\mathbb{E}}_{{\varphi_{t}\left({\bf z}\right)},{\varphi_{t}\left({\bf z}\right)}}[\log p(y_{t}\,|\,{\bf z},{\boldsymbol{\psi}};{\bf x}_{t})]\,,$
${\boldsymbol{\Sigma}}_{t}^{-1}$ $={\boldsymbol{\Sigma}}_{t-1}^{-1}+2\,\nabla_{\boldsymbol{\Sigma}}\,{\mathbb{E}}_{{\phi_{t}\left({\bf w}\right)},{\varphi_{t}\left({\bf z}\right)}}[\log p(y_{t}\,|\,{\bf z},{\bf w};{\bf x}_{t})]\,,$
${\boldsymbol{\Gamma}}_{t}^{-1}$ $={\boldsymbol{\Gamma}}_{t-1}^{-1}+2\,\nabla_{\boldsymbol{\Gamma}}\,{\mathbb{E}}_{{\varphi_{t}\left({\bf z}\right)},{\varphi_{t}\left({\bf z}\right)}}[\log p(y_{t}\,|\,{\bf z},{\boldsymbol{\psi}};{\bf x}_{t})]\,.$
Proof.
First, rewrite (3.6). Let $p(y_{t})\equiv p(y_{t}\,|\,{\bf z},{\bf w};{\bf x}_{t})$ to simplify notation. The loss function is
${\cal K}_{t}$ $=\text{KL}\left({\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}\,{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}\,||\,\phi_{t-1}({\bf w})\,\varphi_{t-1}({\bf z})\,p(y_{t})\right)$
$=\iint{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}\,{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}\log\left(\frac{{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}\,{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}}{{\varphi_{t-1}\left({\bf z}\right)}\,{\phi_{t-1}\left({\bf w}\right)}\,p(y_{t})}\right)\mathrm{d}{\bf z}\,\mathrm{d}{\bf w}$
$=\iint{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}\,{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}\left[\log\left(\frac{{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}}{{\varphi_{t-1}\left({\bf z}\right)}}\right)+\log\left(\frac{{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}}{{\phi_{t-1}\left({\bf w}\right)}}\right)-\log p(y_{t})\right]\mathrm{d}{\bf z}\,\mathrm{d}{\bf w}$
$=\int{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}\,\log\left(\frac{{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}}{{\varphi_{t-1}\left({\bf z}\right)}}\right)\mathrm{d}{\bf z}+\int{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}\log\left(\frac{{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}}{{\phi_{t-1}\left({\bf w}\right)}}\right)\mathrm{d}{\bf w}$
$\quad+\iint{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}\,{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}\,\log p(y_{t})\,\mathrm{d}{\bf w}\,\mathrm{d}{\bf z}$
$=\mathbb{E}_{{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}}\left[\log\left(\frac{{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}}{{\varphi_{t-1}\left({\bf z}\right)}}\right)\right]+\mathbb{E}_{{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}}\left[\log\left(\frac{{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}}{{\phi_{t-1}\left({\bf w}\right)}}\right)\right]$
$\quad+\mathbb{E}_{{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}\,{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}}[\log p(y_{t})]$
$&=\text{KL}\left({\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}\,||\,{\phi_{t-1}\left({\bf w}\right)}\right)+\text{KL}\left({\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}\,||\,{\varphi_{t-1}\left({\bf z}\right)}\right)\\ &\hskip 199.16928pt+\mathbb{E}_{{\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)}\,{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}}[\log p(y_{t})]\,.\end{aligned}$ (A.2)
see Section 6.2.3 in Murphy, (2022).
To simplify notation, let $\mathbb{E}_{{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}{\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}}[\log p(y_{t})]=:{\cal E}_{t}$. The derivative of ${\cal K}_{t}$ with respect to ${\boldsymbol{\nu}}$ is
$\nabla_{\boldsymbol{\nu}}{\cal K}_{t}$ $=\nabla_{\boldsymbol{\nu}}\left(\text{KL}\left({\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)}\,||\,{\phi_{t-1}\left({\bf w}\right)}\right)+{\cal E}_{t}\right)$
$=\nabla_{\boldsymbol{\nu}}\left(\frac{1}{2}{\boldsymbol{\nu}}^{\intercal}{\boldsymbol{\Sigma}}_{t-1}^{-1}{\boldsymbol{\nu}}-{\boldsymbol{\nu}}^{\intercal}{\boldsymbol{\Sigma}}_{t-1}{\boldsymbol{\nu}}_{t-1}+\nabla_{\boldsymbol{\nu}}{\cal E}_{t}\right)$
$={\boldsymbol{\Sigma}}_{t-1}^{-1}{\boldsymbol{\nu}}-{\boldsymbol{\Sigma}}_{t-1}^{-1}{\boldsymbol{\nu}}_{t-1}+\nabla_{\boldsymbol{\nu}}{\cal E}_{t}$
$={\boldsymbol{\Sigma}}_{t-1}^{-1}\left({\boldsymbol{\nu}}-{\boldsymbol{\nu}}_{t-1}-{\boldsymbol{\Sigma}}_{t-1}\nabla_{\boldsymbol{\nu}}{\cal E}_{t}\right).$ (A.3)
Set (A.3) to zero and solve for
${\boldsymbol{\nu}}={\boldsymbol{\nu}}_{t-1}-{\boldsymbol{\Sigma}}_{t-1}\nabla_{\boldsymbol{\nu}}{\cal E}_{t}\,.$
$\nabla_{\boldsymbol{\Sigma}}{\cal K}_{t}$ $=\nabla_{\boldsymbol{\Sigma}}\left(-\frac{1}{2}\,\log|{\boldsymbol{\Sigma}}|+\frac{1}{2}\,\text{Tr}\left({\boldsymbol{\Sigma}}\,{\boldsymbol{\Sigma}}_{t-1}^{-1}\right)+{\cal E}_{t}\right)$
$=-\frac{1}{2}\,{\boldsymbol{\Sigma}}^{-1}+\frac{1}{2}\,{\boldsymbol{\Sigma}}_{t-1}^{-1}+\nabla_{{\boldsymbol{\Sigma}}}{\cal E}_{t}.$ (A.4)
The fixed-point solution for (A.4) satisfies
${\boldsymbol{\Sigma}}^{-1}={\boldsymbol{\Sigma}}_{t-1}^{-1}+2\,\nabla_{{\boldsymbol{\Sigma}}}{\cal E}_{t}\,.$
Corollary 1.
Suppose $\log p(y\,|\,{\bf z},{\bf w})$ is differentiable with respect to $({\bf z},{\bf w})$ and the observations $\{y_{t}\}_{t=1}^{T}$ are conditionally independent over $({\bf z},{\bf w})$. Given Gaussian prior distributions $\phi_{0}$, $\varphi_{0}$ for ${\bf w}$ and ${\bf z}$ respectively, the modified R-VGA equations for PULSE in order-2 form are
${\boldsymbol{\nu}}_{t}$ $={\boldsymbol{\nu}}_{t-1}+{\boldsymbol{\Sigma}}_{t-1}{\mathbb{E}}_{{\phi_{t}\left({\bf w}\right)}\,{\varphi_{t}\left({\bf z}\right)}}[\nabla_{\bf w}\log p(y_{t}\,|\,{\bf z},{\bf w};{\bf x}_{t})]\,,$
${\boldsymbol{\mu}}_{t}$ $={\boldsymbol{\mu}}_{t-1}+{\boldsymbol{\Gamma}}_{t-1}{\mathbb{E}}_{{\phi_{t}\left({\bf w}\right)}\,{\varphi_{t}\left({\bf z}\right)}}[\nabla_{\bf z}\log p(y_{t}\,|\,{\bf z},{\bf w};{\bf x}_{t})]\,,$
${\boldsymbol{\Sigma}}_{t}^{-1}$ $={\boldsymbol{\Sigma}}_{t-1}^{-1}-{\mathbb{E}}_{{\phi_{t}\left({\bf w}\right)}\,{\varphi_{t}\left({\bf z}\right)}}[\nabla_{\bf w}^{2}\log p(y_{t}\,|\,{\bf z},{\bf w};{\bf x}_{t})]\,,$
${\boldsymbol{\Gamma}}_{t}^{-1}$ $={\boldsymbol{\Gamma}}_{t-1}^{-1}-{\mathbb{E}}_{{\varphi_{t}\left({\bf z}\right)}\,{\varphi_{t}\left({\bf z}\right)}}[\nabla_{\bf z}^{2}\log p(y_{t}\,|\,{\bf z},{\bf w};{\bf x}_{t})]\,.$
Proof.
The proof follows directly by rearranging the order of integration, Bonnet’s Theorem, and Prices’s Lemma, see Lemma 1 and Theorem 1 in Lin et al., (2019). ∎
Proposition 2.
Given a logistic regression model written in the canonical form
$\log p(y_{t})$ $=y_{t}\log\left(\frac{\sigma({\bm{\theta}}^{\intercal}\,{\bf x}_{t})}{1-\sigma(\bm{\theta}^{\intercal}\,{\bf x}_{t})}\right)-\left(1+\exp(\sigma(\bm{\theta}^{\intercal}\,{\bf x}_{t}))\right)\,,$
we have that the gradient of $\log\left(\frac{\sigma({\bm{\theta}}^{\intercal}\,{\bf x}_{t})}{1-\sigma(\bm{\theta}^{\intercal}\,{\bf x}_{t})}\right)$ with respect to $\bm{\theta}$ is given by
$\nabla_{\bm{\theta}}\log\left(\frac{\sigma({\bm{\theta}}^{\intercal}\,{\bf x}_{t})}{1-\sigma(\bm{\theta}^{\intercal}\,{\bf x}_{t})}\right)={\bf x}_{t}\,.$
Proof.
$\nabla_{\bm{\theta}}\log\left(\frac{\sigma({\bm{\theta}}^{\intercal}\,{\bf x}_{t})}{1-\sigma(\bm{\theta}^{\intercal}\,{\bf x}_{t})}\right)$ $=\nabla_{\bm{\theta}}\log\left(\sigma({\bm{\theta}}^{\intercal}\,{\bf x}_{t}\right)-\nabla_{\bm{\theta}}\log\left(1-\sigma(\bm{\theta}^{\intercal}\,{\bf x}_{t})\right)$
$=\left(\sigma({\bm{\theta}}^{\intercal}\,{\bf x}_{t})\right)^{-1}\sigma^{\prime}({\bm{\theta}}^{\intercal}\,{\bf x}_{t})\,{\bf x}_{t}+\left(1-\sigma(\bm{\theta}^{\intercal}\,{\bf x}_{t})\right)^{-1}\sigma^{\prime}({\bm{\theta}}^{\intercal}\,{\bf x}_{t}){\bf x}_{t}$
$=(1-\sigma(\bm{\theta}^{\intercal}\,{\bf x}_{t})+\sigma(\bm{\theta}^{\intercal}\,{\bf x}_{t}))\,{\bf x}_{t}$
$={\bf x}_{t}\,.$
A.1 Proof of Theorem 2
First rewrite the log-likelihood of the target variable $y_{t}$ as a member of the exponential-family. Let $f_{t}({\bf z},{\bf w})={\bf w}^{\intercal}\,h({\bf z};{\bf x}_{t})$. Then,
$\log p(y_{t})$ $=\log\text{Bern}\Big{(}y_{t}\,|\,\sigma({\bf w}^{\intercal}\,h({\bf z};{\bf x}_{t}))\Big{)}$
$=y_{t}\log\sigma(f_{t}({\bf z},{\bf w}))+(1-y_{t})\log(1-\sigma(f_{t}({\bf z},{\bf w})))$
$=y_{t}\log\left(\frac{\sigma(f_{t}({\bf z},{\bf w}))}{1-\sigma(f_{t}({\bf z},{\bf w}))}\right)+\log(1-\sigma(f_{t}({\bf z},{\bf w})))$
$=y_{t}\log\left(\frac{\sigma(f_{t}({\bf z},{\bf w}))}{1-\sigma(f_{t}({\bf z},{\bf w}))}\right)-\log\left(1+\exp\left(\log\left(\frac{\sigma(f_{t}({\bf z},{\bf w}))}{1-\sigma(f_{t}({\bf z},{\bf w}))}\right)\right)\right)$
$=y_{t}\,\eta_{t}-\log(1+\exp(\eta_{t}))$
$=y_{t}\,\eta_{t}-A(\eta_{t})\,,$
where $\eta_{t}=\log\left(\frac{\sigma(f_{t}({\bf z},{\bf w}))}{1-\sigma(f_{t}({\bf z},{\bf w}))}\right)$ is the natural parameter and $A(\eta_{t})$ the log-partition function. We perform a moment-match estimation. To do this, we follow the property of exponential-family distributions, and use the results in Proposition 2. The first and second-order derivatives of the log-partition $A(\eta_{t})$ are
$\frac{\partial}{\partial\eta_{t}}A(\eta_{t})$ $={\mathbb{E}}[y\,|\,\eta_{t}]=\sigma(f_{t}({\bf z},{\bf w}))\,,$ (A.5)
$\frac{\partial^{2}}{\partial\eta_{t}^{2}}A(\eta_{t})$ $=\text{Cov}[y\,|\,\eta_{t}]=\sigma^{\prime}(f_{t}({\bf z},{\bf w}))\,.$ (A.6)
Then, the first order approximation $\hat{\sigma}(f_{t}({\bf z},{\bf w}))$ is
$\hat{\sigma}(f_{t}({\bf z},{\bf w}))$ $=\sigma(\bar{f}_{t-1})+\sigma^{\prime}(\bar{f}_{t-1})(F_{t,{\bf z}}^{\intercal}\,({\bf z}-{\boldsymbol{\mu}}_{t-1})+F_{t,{\bf w}}^{\intercal}\,({\bf w}-{\boldsymbol{\nu}}_{t-1}))\,,$
where $\bar{f}_{t-1}=f_{t}({\boldsymbol{\mu}}_{t-1},{\boldsymbol{\nu}}_{t-1})$, $F_{t,{\bf z}}=\nabla_{{\bf z}}f_{t}({\bf z},{\boldsymbol{\nu}}_{t-1})\big{|}_{{\bf z}={\boldsymbol{\mu}}_{t-1}}$, and $F_{t,{\bf w}}=\nabla_{{\bf w}}f_{t}({\boldsymbol{\mu}}_{t-1},{\bf w})\big{|}_{{\bf w}={\boldsymbol{\nu}}_{t-1}}=h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t})$. Next, we derive the update equations for ${\boldsymbol{\nu}}$ and ${\boldsymbol{\Sigma}}$. The moment-matched log-likelihood is given by
$\log p(y_{t})=\log\mathcal{N}\left({y_{t}}\,|\,{\hat{\sigma}(f_{t}({\bf z},{\bf w}))},\,{\sigma(f_{t}({\bf z},{\bf w}))(1-\sigma(f_{t}({\bf z},{\bf w})))}\right)$ (A.7)
and the gradient with respect to ${\bf w}$ is
$\nabla_{\bf w}\log p(y_{t})=y_{t}\,F_{t,{\bf w}}-\hat{\sigma}(\bar{f}_{t-1})\,F_{t,{\bf w}}=(y_{t}-\hat{\sigma}(\bar{f}_{t-1}))\,F_{t,{\bf w}}\,.$
Now, the Hessian of the log-model with respect to ${\bf w}$ is
$\nabla_{\bf w}^{2}\log p(y_{t})$ $=\nabla_{\bf w}(y_{t}-\hat{\sigma}(\bar{f}_{t-1}))\,F_{t,{\bf w}}$
$=-\sigma^{\prime}(\bar{f}_{t-1})\,F_{t,{\bf w}}^{\intercal}\,\,F_{t,{\bf w}}\,.$
The update for ${\boldsymbol{\Sigma}}_{t}$, following the order-2 form of the modified equations and replacing the expectation under ${\varphi_{t}\left({\bf z}\right)}$ for ${\varphi_{t-1}\left({\bf z}\right)}$, is
${\boldsymbol{\Sigma}}_{t}^{-1}$ $={\boldsymbol{\Sigma}}_{t-1}^{-1}-{\mathbb{E}}_{{\phi_{t}\left({\bf w}\right)}\,{\varphi_{t-1}\left({\bf z}\right)}}\left[\nabla^{2}_{\bf w}\log p(y_{t})\right]$
$={\boldsymbol{\Sigma}}_{t-1}^{-1}-{\mathbb{E}}_{{\phi_{t}\left({\bf w}\right)}\,{\varphi_{t-1}\left({\bf z}\right)}}\left[-\sigma^{\prime}(\bar{f}_{t-1})\,F_{t,{\bf w}}^{\intercal}\,F_{t,{\bf w}}\right]$
$={\boldsymbol{\Sigma}}_{t-1}^{-1}+\sigma^{\prime}(\bar{f}_{t-1})\,F_{t,{\bf w}}^{\intercal}\,F_{t,{\bf w}}$
$={\boldsymbol{\Sigma}}_{t-1}^{-1}+\sigma^{\prime}\big{(}{\boldsymbol{\nu}}_{t-1}^{\intercal}\,h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t})\big{)}\,h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t})^{\intercal}\,h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t})\,.$
Next, the update step for ${\boldsymbol{\nu}}_{t}$ becomes
${\boldsymbol{\nu}}_{t}$ $={\boldsymbol{\nu}}_{t-1}-{\boldsymbol{\Sigma}}_{t-1}\,{\mathbb{E}}_{{\phi_{t}\left({\bf w}\right)}{\varphi_{t-1}\left({\bf z}\right)}}[\nabla_{\bf w}\log p(y_{t}\,|\,{\bf z},{\bf w};{\bf x}_{t})]$
$={\boldsymbol{\nu}}_{t-1}+{\boldsymbol{\Sigma}}_{t-1}\,{\mathbb{E}}_{{\phi_{t}\left({\bf w}\right)}{\varphi_{t-1}\left({\bf z}\right)}}[(y_{t}-\hat{\sigma}(\bar{f}_{t-1}))\,F_{t,{\bf z}}]$
$={\boldsymbol{\nu}}_{t-1}+{\boldsymbol{\Sigma}}_{t-1}\,{\mathbb{E}}_{{\phi_{t}\left({\bf w}\right)}{\varphi_{t-1}\left({\bf z}\right)}}\Big{[}(y_{t}-\sigma(\bar{f}_{t-1})$
$\qquad+\sigma^{\prime}(\bar{f}_{t-1})(F_{t,{\bf z}}({\bf z}-{\boldsymbol{\mu}}_{t-1})+F_{t,{\bf w}}^{\intercal}\,({\bf w}-{\boldsymbol{\nu}}_{t-1}))\Big{]}F_{t,{\bf z}}$
$={\boldsymbol{\nu}}_{t-1}+{\boldsymbol{\Sigma}}_{t-1}\,F_{t,{\bf w}}\Big{(}y_{t}-\sigma(\bar{f}_{t-1})-\sigma^{\prime}(\bar{f}_{t-1})\,F_{t,{\bf w}}^{\intercal}\,({\boldsymbol{\nu}}_{t}-{\boldsymbol{\nu}}_{t-1})\Big{)}$
$={\boldsymbol{\nu}}_{t-1}+{\boldsymbol{\Sigma}}_{t-1}\,F_{t,{\bf w}}\Big{(}y_{t}-\sigma(\bar{f}_{t-1})+\sigma^{\prime}(\bar{f}_{t-1})\,F_{t,{\bf w}}^{\intercal}\,{\boldsymbol{\nu}}_{t-1})\Big{)}-\sigma^{\prime}(\bar{f}_{t-1}){\boldsymbol{\Sigma}}_{t-1}\,F_{t,{\bf w}}\,F_{t,{\bf w}}^{\intercal}\,{\boldsymbol{\nu}}_{t}\,.$
We rewrite the last equality as
${\boldsymbol{\nu}}_{t}+\sigma^{\prime}(\bar{f}_{t-1}){\boldsymbol{\Sigma}}_{t-1}\,F_{t,{\bf w}}\,F_{t,{\bf w}}^{\intercal}\,{\boldsymbol{\nu}}_{t}={\boldsymbol{\nu}}_{t-1}+{\boldsymbol{\Sigma}}_{t-1}\,F_{t,{\bf w}}\Big{(}y_{t}-\sigma(\bar{f}_{t-1})+\sigma^{\prime}(\bar{f}_{t-1})\,F_{t,{\bf w}}^{\intercal}\,{\boldsymbol{\nu}}_{t-1})\Big{)}\,,$
which implies that
$\left({\bf I}+\sigma^{\prime}(\bar{f}_{t-1})\,{\boldsymbol{\Sigma}}_{t-1}\,F_{t,{\bf w}}\,F_{t,{\bf w}}^{\intercal}\,\right){\boldsymbol{\nu}}_{t}={\boldsymbol{\nu}}_{t-1}+{\boldsymbol{\Sigma}}_{t-1}\,F_{t,{\bf w}}\,\Big{(}y_{t}-\sigma(\bar{f}_{t-1})+\sigma^{\prime}(\bar{f}_{t-1})\,F_{t,{\bf w}}^{\intercal}\,{\boldsymbol{\nu}}_{t-1})\Big{)}\,.$
Similarly, we have that
${\boldsymbol{\nu}}_{t}=\left({\bf I}+\sigma^{\prime}(\bar{f}_{t-1}){\boldsymbol{\Sigma}}_{t-1}\,F_{t,{\bf w}}F_{t,{\bf w}}^{\intercal}\,\right)^{-1}\left({\boldsymbol{\nu}}_{t-1}+{\boldsymbol{\Sigma}}_{t-1}F_{t,{\bf w}}\Big{(}y_{t}-\sigma(\bar{f}_{t-1})+\sigma^{\prime}(\bar{f}_{t-1})\,F_{t,{\bf w}}^{\intercal}\,{\boldsymbol{\nu}}_{t-1})\Big{)}\right)\,,$
where
$\left({\bf I}+\sigma^{\prime}(\bar{f}_{t-1})\,{\boldsymbol{\Sigma}}_{t-1}\,F_{t,{\bf w}}F_{t,{\bf w}}^{\intercal}\,\right)^{-1}=$ $\left({\boldsymbol{\Sigma}}_{t-1}\left[{\boldsymbol{\Sigma}}_{t-1}^{-1}+\sigma^{\prime}(\bar{f}_{t-1})F_{t,{\bf w}}F_{t,{\bf w}}^{\intercal}\,\right]\right)^{-1}$
$=$ $\left[{\boldsymbol{\Sigma}}_{t-1}^{-1}+\sigma^{\prime}(\bar{f}_{t-1})F_{t,{\bf w}}F_{t,{\bf w}}^{\intercal}\,\right]^{-1}{\boldsymbol{\Sigma}}_{t-1}^{-1}$
$=$ $\,{\boldsymbol{\Sigma}}_{t}\,{\boldsymbol{\Sigma}}_{t-1}^{-1}\,.$
Then, it follows that
${\boldsymbol{\nu}}_{t}$ $={\boldsymbol{\Sigma}}_{t}\,{\boldsymbol{\Sigma}}_{t-1}^{-1}\left({\boldsymbol{\nu}}_{t-1}+{\boldsymbol{\Sigma}}_{t-1}\,F_{t,{\bf w}}\Big{(}y_{t}-\sigma(\bar{f}_{t-1})+\sigma^{\prime}(\bar{f}_{t-1})\,F_{t,{\bf w}}^{\intercal}\,{\boldsymbol{\nu}}_{t-1})\Big{)}\right)$
[…]

### v2

@@ C.3 The added value of one model per client
For PULSE, with the exception of Client 1, the performance of one model per client is significantly lower than that of the universal model. Indeed, the additional data are more valuable than a model per client. The universal model employs features that are built using the identity of the client, e.g., inventory and cash of client. Thus, it is more advantageous to have one model for all clients and benefit from more data points than one model per client at the expense of having fewer data points to train the individual models. This is the case for PULSE and RF, whereas the results for LogR and MLE are not conclusive.
The above results are for a universal model that does employ client-specific features. If we consider a universal model without access to client-specific features (e.g., cash, inventory, and recent activity) the performance of the models deteriorates. For example, for Client 1 (using PULSE) we find close to 20% decrease in the ability to predict toxic flow when going from a universal model with client-specific features to a universal model without client-specific features.
C.4 Global models
In this section we study the added value of employing data from more clients to build the models. Figure 19 reports the median daily AUC across toxicity horizons under two training regimes: (i) a single global model and (ii) client-specific models (one model per client). The global model is fit on pooled data from the top 100 clients and excludes client-identifying features. The client-specific models are trained separately for the top six clients and include client information features.
Figure 19: Each panel shows daily AUC versus toxicity horizon for one method under two configurations: (i) per-client models and (ii) a single global model. Solid lines denote the global configuration; dashed lines denote per-client; the bands around each line indicate the interquartile range (IQR). EUR/USD, 1 Aug 2022–21 Oct 2022.
Pooling data helps flexible models. The mean daily AUC is higher for PULSE and RF when using a single global model than when fitting one model per client—these methods can exploit the larger sample and share statistical strength across clients. LogR performs roughly on par in both settings, though the per-client variant shows higher variance due to smaller sample sizes. MLE is the exception: a per-client MLE (client-specific base rate) outperforms the global MLE, since a global average dilutes client idiosyncrasies. Overall, PULSE is the top performer, followed by RF for both training regimes. Thus, in our dataset, having more clients (more data) is advantageous. This finding is specific to our dataset (it is not universal); this might be a consequence of the pool of clients we study, the features we consider, or the period in our analysis.
C.5 Weekly re-trains
In this section, we compare the performance of the NNet trained with PULSE to the performance of the RF and LogR models with weekly re-training. At the beginning of each week of the deploy phase, we use the previous week’s data to re-train RF and LogR. For the first week in the deploy phase, we make predictions with data of the warmup phase. Figure 20 shows the five-day rolling AUC mean with $\mathcal{G}=10s$.
Figure 20: Five-day exponentially-weighted moving average of AUC over time. The toxicity horizon is ten seconds.
The decrease in performance for RF and LogR with weekly retrain is less pronounced than when training only in the deploy phase. As above, the mean AUC of RF and LogR with weekly re-retraining is lower than that of the NNet trained with PULSE.
The NNet trained with PULSE only observes each datapoint once, whereas RF keeps the data in memory to re-train. Thus, a NNet trained with PULSE is more memory-efficient and more scalable than both RF and LogR. Finally, given that PULSE is fully-online, it reacts more quickly to changes in the behaviour of clients.
C.6 Variance of LogR
Here, we investigate the variance we observed in Figure 9, Figure 17, and the concentrated probabilities shown in Figure 13.
Figure 21 shows the normalised coefficient magnitudes $|w_{j}|/\sum_{k}|w_{k}|$ for per-client LogR fits. We observe a marked concentration of mass: few coefficients (in some instances, one) dominate, while the remainder are near zero, and the index of the dominant coefficient varies across clients. This reflects weak linear signal and collinearity in small per-client samples. Consistent with this, LogR’s predicted probabilities concentrate near a base rate, yielding near-step internalised-volume curves and the horizon-to-horizon jaggedness noted in the main text.
Figure 21: Normalised coefficient magnitudes $|w_{d}|/\sum_{j}|w_{j}|$ for per-client LogR fits (top six clients).
C.7 Precision and recall
[…]